a strong current of trivalent gaseous boron passed through germanium crystal decreases the density of the crystal due to part replacement of germanium by boron and due to interstitial vacancies is created by missing germanium atoms. in one such experiment one gram germanium is taken an d boron atoms are found to be 150 ppm by weight when the density of germanium crystal decreases by 4%. calculate the percentage of missing vacancies due to germanium which are filled up by boron atoms.
1
Expert's answer
2015-06-05T12:26:26-0400
1 g Germanium------- 150 ppm = 1.5 * 10^-4 g Boron

4% decrease---- now weight is 0.96 g(Ge + B)----0.04 g missing

0.04 missing-------- (0.04-0.00015)g Ge missing------(0.04-0.00015)/72.61 moles
Ge missing=5.488*10^-4 moles

Boron replaced= 1.5*10^-4 g = (1.5*10^-4)/10 moles

equivalently replaced = 3/4((1.5*10^-4)/10) moles of Ge ( Valency of B is 3 and
Ge is 4)

= 0.1125*10^-4 moles of Ge

% of Missing vacancies filled by B atoms = 0.1125*100/5.488 = 2.05 %